Equations of quadratic form
4x to the 4th power-10xsquare plus 9=0 I'm not sure whata numbers to factor

Answer by stanbon(75887)

4x to the 4th power-10xsquare plus 9=0 I'm not sure whata numbers to factor
-----
4x^4 - 10x^2 + 9 = 0
----
Use the Quadratic Formula:
x^2 = [10 +- sqrt(100-4*4*9)]/(2*4)
---
x^2 = [10 +- sqrt(-17)]/8
---
Solve for "x":
=====
Cheers,
Stan H.
======